# spacing=nonfrench

### 20 Mar 2026 v1

@@ 4.1 Isabelle REPL
To enable efficient interaction with Isabelle from Python, we implement a new Isabelle REPL (Read-Eval-Print Loop), [^2] Isabelle REPL is an interactive command-line interface for the Isabelle theorem prover, designed to support incremental, exploratory proof development, which is similar in spirit to the REPLs found in programming languages such as Python, but adapted to Isabelle’s document-oriented proof model. whose logic builds on scala-isabelle unruh2025scalaisabelle and Py4J (nishant2023py4j, ). The REPL provides a Gateway Server that exposes Isabelle components to Python clients through a unified API. This design enables programmatic interaction with Isabelle, allowing users to parse Isabelle theories, extract run-time context information, and execute proof steps incrementally.
@@ 4.2 Experimental Setup
Dataset. We use the FVELER dataset (lin2024fvel, ), which comprises 29,125 theorems from seL4. The dataset is split into four subsets, i.e., training, validation, test, and test-hard, containing 26,081, 1,115, 1,077, and 852 theorems, respectively. Following the FVEL criteria, the training, validation, and test sets are randomly partitioned. In contrast, the test-hard set consists of theorems drawn from three specific and independent sessions, i.e., SysInit, SysInitExamples, and LibTest, that do not appear in any of the other splits. Note that sessions are the way Isabelle organizes theory files, which resemble the relationship between code files and libraries in other programming languages. A theory can import other theories, and a session can depend on other sessions by declaring dependencies in the Isabelle ROOT file. These sessions are typically selected based on the depth of their session-level dependency graphs, and the theorems they contain exhibit more complex dependencies, making them considerably more challenging to prove.
@@
Training. We fine-tuned two small models, Qwen3-1.7B (yang2025qwen3, ) and Mistral-7B (jiang2023mistral, ), to predict and score each proof step. We carried out the full-parameter supervised fine-tuning (SFT) using the Llama-Factory framework (zheng2024llamafactory, ). We use DeepSpeed ZeRO-2 optimization to enable memory-efficient distributed training. The training used an effective global batch size of 16 and ran 3 epochs using bfloat16 mixed precision. The learning rate was set to 1e-5 and followed a cosine decay schedule with a warmup ratio of 0.1.
@@
Baseline. We select four neural or symbolic proof-generation methods as baselines. Selene prompts GPT-4o with the target problem and several examples to generate a candidate proof. FVEL follows a similar paradigm but also fine-tunes a Mistral 7B-Instruct model on the training data to produce complete proofs, rather than using in-context learning. We also include two symbolic methods, Auto and Sledgehammer, which operate purely symbolically to prove the target theorem. Here, Auto is a collection of common symbolic automatic proving methods in Isabelle, such as by simp, by auto, and by fastforce. The time limit for all methods is set to 120 minutes, as we observed no noticeable improvement when increasing it further.
@@ 4.3.1 RQ1: Effectiveness Comparison
Table 1: Proof success rates (%) across validation (Val), test (Test), and test-hard (TsHd) splits. Compared to existing neural and symbolic provers, our framework delivers substantial improvements: Qwen3-1.7B achieves 70.4% overall success, and Mistral-7B reaches 77.6%, exceeding the strongest baseline by 30.0% and 37.3%, respectively.
@@
Table 1 summarizes the performance of our approach compared to both neural and symbolic baselines. Among neural methods, Selene achieved 156 successful proofs, while FVEL improved this to 219. As for symbolic baselines, Auto proved 164 theorems, whereas Sledgehammer reached 1,124. In contrast, our framework substantially outperforms all these baselines across different splits. Using the fine-tuned Mistral-7B model, our approach produces a total of 2,167 feasible proofs, of which 788 are on the validation set, 811 are on the test set, and 568 are on the test-hard set. This accounts for 77.6% of all still valid theorems across the three evaluated sets and proves 37.3% more theorems than the best baseline, Sledgehammer. It is also encouraging that, on a relatively small model, Qwen3-1.7B, our approach manages to complete the proof of 1,965 theorems, which amounts to 70.4% of the total theorems. This demonstrates that in a real environment where new proofs need to be crafted, a deployed small model can already help with the development of verification.
We further evaluate the performance of our automated proving approach in two aspects. First, prior work has struggled with generating nontrivial proofs. Since the proofs in seL4 are largely procedural, we use the proof length as a proxy for proving difficulty. We compute the proof success rate across different proof lengths and present the results in Figure 3. The results show that our approach is more capable of proving longer theorems than previous baselines. Although the success rate naturally decreases as proof length increases, it remains around 20% even for the 393 theorems whose proofs exceed 10 lines. We leave it as future work to investigate more effective strategies for constructing longer proofs.
@@
Response to RQ1 Our approach substantially outperforms baselines, producing 37.3% more proofs than the (symbolic) ATP tool Sledgehammer and 69.8% more proofs than the state-of-the-art neural method. It remains reliably effective across different proof difficulty levels and session categories. Further, it demonstrates good generalizability, successfully discharging 67.6% of the theorems from unseen sessions during training.
@@ 4.3.2 RQ2: Proving Efficiency in Saving Human Efforts
We further situate our method within an AI–human collaborative pipeline to evaluate its efficiency in reducing human effort. Specifically, for each theorem, we provide a portion of the ground-truth proof lines (simulating human expert-written proof) and ask our approach to complete the remaining steps. Figure 5 shows the proof completion rate curve w.r.t. the proportion of ground-truth proof.
As expected, supplying more expert-written proof lines enables our method to complete more theorems. Ultimately, including cases solved entirely automatically, our approach succeeds in automatically completing at least 10% export effort for 86.0% of the theorems. We then calculate the average effort saving (AES) of our approach:
$\text{AES}=p_{1}\sigma_{1}+\sum_{i=2}^{k}(p_{k}-p_{k-1})\sigma_{k},$
where $(\sigma_{1},\dots,\sigma_{k})$ are the ratios of expert-written proof lines in ascending order, with the corresponding proof completion rates are $(p_{1},\dots,p_{k})$. Therefore, under this AI–human collaboration setting, the automated proof completion reduces human effort by an average of 71.1%.
We also record the time consumption of our framework as a reference; the average time to generate a successful proof is 139.1 minutes. However, it is worth noting that this average is skewed upward by a small number of theorems that take extremely long to prove. In our experiment, we find that 58.4% of the theorems are proved within 10 minutes, 73.4% within 30 minutes, and 80.8% within 2 hours. Therefore, the majority of proofs are obtained within a relatively short time horizon, indicating that our framework is generally highly efficient despite a few outliers with long proof times.
@@ 4.3.3 RQ3: Generalizability to Other Projects
A major concern with training-based methods is their generazability, namely, whether the approach is only applicable to a specific domain due to potential data leakage or overfitting.
To investigate possible leakage, given that the seL4 proofs are publicly available and that the pre-trained model may have already incorporated some of the knowledge, we compute the similarity between the ground-truth proofs and the proofs generated by our approach. We adopt sequence similarity (lcvenshtcin1966binary, ) and Jaccard similarity (jaccard1901etude, ) as the evaluation metrics. The results are shown in Figure 6. They indicate that both similarity scores remain low except for one- or two-line proofs, which is not surprising since very short proofs tend to be identical. We further examined these short proofs and found that only 16.7% are identical, providing strong evidence that data leakage in our fine-tuned model is unlikely.
Figure 6: Sequence and Jaccard similarity between generated and ground-truth proofs. (Note that sequence similarity measures how alike two ordered sequences are based on their matching elements and positions, while Jaccard similarity measures how similar two sets are by comparing the proportion of shared elements to the total distinct elements.) Both similarities stay below the thresholds of indicating potential data leakage across different proof lengths.
We further explore the generalizability of our approach and the fine-tuned LLM. We select three projects from the Archive of Formal Proofs (AFP [^3] http://afp.sf.org), i.e., X86 Semantics X86_Semantics-AFP , IEEE Floating Point IEEE_Floating_Point-AFP , and SATSolverVerification SATSolverVerification-AFP , to directly apply proof search on the theorems therein. In addition, we include the code verification benchmark Code2Inv and translate its loop invariants into Isabelle theorems.
@@ 4.3.4 RQ4: Ablation study
Since our approach integrates several components, namely LLMs, hammer, as well as proof-step revision and proof-state filtering mechanisms, we conduct ablation studies to evaluate the contribution of each part of the framework. To assess the impact of our fine-tuned model, we replace it with off-the-shelf LLMs such as GPT-4o or DeepSeek. Due to budget constraints (running tree search for a single problem costs approximately \$4.1 when using the DeepSeek or GPT-4o APIs), we randomly sample 200 theorems from the benchmark and use the DeepSeek API to propose proof steps.
In addition, inspired by Rango (thompson2024rango, ), we implement a retrieval-augmented generation (RAG) mechanism. The retrieval module fetches both similar proof steps and potentially relevant premises. Similar steps are retrieved from the training set based on the similarity between their associated proof states and the current proof state, where these states are tokenized into sequences of identifiers, and relevance is measured using BM25 (robertson2009probabilistic, ). For premises retrieval, we continue to rely on Sledgehammer’s MePo heuristic for efficiency. The final prompt for LLMs includes the retrieved steps and premises along with their statements, and we restrict the total lengths of steps and premises to 1024 and 512 tokens, respectively.
@@
Regarding the necessity of proof state filtering, particularly the counterexample detection, our experiments show that incorporating this component allows the system to identify an average of 1.3 counterexamples per evaluated theorem. For the subset of search traces that contain counterexamples, the average number rises to 11.1. These results indicate that counterexample detection is valuable for identifying unprovable states early and preventing unproductive search efforts, although it does not directly increase the overall proof success rate.

### 07 May 2026 v2

@@ 4.1 Isabelle REPL
To enable efficient interaction with Isabelle from Python, we implement a new Isabelle REPL (Read-Eval-Print Loop), whose logic builds on scala-isabelle unruh2025scalaisabelle and Py4J (nishant2023py4j). The REPL provides a Gateway Server that exposes Isabelle components to Python clients through a unified API. This design enables programmatic interaction with Isabelle, allowing users to parse Isabelle theories, extract run-time context information, and execute proof steps incrementally.
@@ 4.2 Experimental Setup
Dataset. We use the FVELER dataset (lin2024fvel), which comprises 29,125 theorems from seL4. The dataset is split into four subsets, i.e., training, validation, test, and test-hard, containing 26,081, 1,115, 1,077, and 852 theorems, respectively. Following the FVEL criteria, the training, validation, and test sets are randomly partitioned. In contrast, the test-hard set consists of theorems drawn from three specific and independent sessions, i.e., SysInit, SysInitExamples, and LibTest, that do not appear in any of the other splits. Note that sessions are the way Isabelle organizes theory files, which resemble the relationship between code files and libraries in other programming languages. A theory can import other theories, and a session can depend on other sessions by declaring dependencies in the Isabelle ROOT file. These sessions are typically selected based on the depth of their session-level dependency graphs, and the theorems they contain exhibit more complex dependencies, making them considerably more challenging to prove.
@@
Training. We fine-tuned two small models, Qwen3-1.7B (yang2025qwen3) and Mistral-7B (jiang2023mistral), to predict and score each proof step. We carried out the full-parameter supervised fine-tuning (SFT) using the Llama-Factory framework (zheng2024llamafactory). We use DeepSpeed ZeRO-2 optimization to enable memory-efficient distributed training. The training used an effective global batch size of 16 and ran 3 epochs using bfloat16 mixed precision. The learning rate was set to 1e-5 and followed a cosine decay schedule with a warmup ratio of 0.1.
@@
Baseline. We select four neural or symbolic proof-generation methods as baselines. Selene prompts GPT-4o with the target problem and several examples to generate a candidate proof. FVEL follows a similar paradigm but also fine-tunes a Mistral-7B-Instruct model on the training data to produce complete proofs, rather than using in-context learning. We also include two symbolic methods, Auto and Sledgehammer, which operate purely symbolically to prove the target theorem. Here, Auto is a collection of common symbolic automatic proving methods in Isabelle, such as by simp, by auto, and by fastforce. The time limit for all methods is set to 120 minutes, as we observed no noticeable improvement when increasing it further.
@@ 4.3.1 RQ1: Effectiveness Comparison
Table 1: Proof success rates (%) across validation (Val), test (Test), and test-hard (TsHd) splits. Compared to existing neural and symbolic provers, our framework delivers substantial improvements: Qwen3-1.7B achieves 70.4% overall success, and Mistral-7B reaches 77.6%, exceeding the strongest baseline by 30.0 and 37.3 percentage points, respectively.
@@
Table 1 summarizes the performance of our approach compared to both neural and symbolic baselines. Among neural methods, Selene achieved 156 successful proofs, while FVEL improved this to 219. As for symbolic baselines, Auto proved 164 theorems, whereas Sledgehammer reached 1,124. In contrast, our framework substantially outperforms all these baselines across different splits. Using the fine-tuned Mistral-7B model, our approach produces a total of 2,167 successful proofs, of which 788 are on the validation set, 811 are on the test set, and 568 are on the test-hard set. This accounts for 77.6% of all still valid theorems across the three evaluated sets, 37.3 percentage points higher than the best baseline, Sledgehammer. It is also encouraging that, on a relatively small model, Qwen3-1.7B, our approach manages to complete the proof of 1,965 theorems, which amounts to 70.4% of the total theorems. This demonstrates that in a real environment where new proofs need to be crafted, a deployed small model can already help with the development of verification.
We further evaluate the performance of our automated proving approach in two aspects. First, prior work struggled to generate nontrivial proofs. Since the proofs in seL4 are largely procedural, we use the proof length as a proxy for proving difficulty. We compute the proof success rate across different proof lengths and present the results in Figure 3. The results show that our approach is more capable of proving longer theorems than previous baselines. Although the success rate naturally decreases as proof length increases, it remains around 20% even for the 393 theorems whose proofs exceed 10 lines. We leave it as future work to investigate more effective strategies for constructing longer proofs.
@@
Response to RQ1 Our approach substantially outperforms baselines, achieving a success rate 37.3 percentage points higher than the (symbolic) ATP tool Sledgehammer and 69.8 percentage points higher than the state-of-the-art neural method. It remains reliably effective across different proof difficulty levels and session categories. Further, it demonstrates good generalizability, successfully discharging 67.6% of the theorems from unseen sessions during training.
@@ 4.3.2 RQ2: Proving Efficiency in Saving Human Efforts
To evaluate how effectively our approach reduces manual proof effort, we simulate an AI–human collaborative workflow: for each theorem, we supply a prefix of the ground-truth proof (the first $\sigma$ fraction of expert-written lines) and ask our approach to complete the remaining steps. Figure 5 plots the proof completion rate as a function of the prefix ratio $\sigma$.
As the prefix grows, more theorems become completable. Including cases solved entirely automatically ($\sigma=0$), our approach saves at least 10% of the expert proof effort for 86.0% of the theorems. To summarize this curve with a single metric, we define the average effort saving $\eta$:
$\eta=\sum_{i=1}^{k}(p_{i}-p_{i-1})\cdot\sigma_{i},\quad p_{0}=0,$
where $(\sigma_{1},\dots,\sigma_{k})$ are the prefix ratios in ascending order, and $p_{i}$ is the fraction of theorems whose remaining proof can be completed automatically given a $\sigma_{i}$-prefix. This yields $\eta=71.1\%$, meaning our approach reduces human proof effort by roughly two-thirds on average.
We also record the time consumption of our framework as a reference; the average time to generate a successful proof is 139.1 minutes. However, it is worth noting that this average is skewed upward by a small number of theorems that take extremely long to prove. In our experiment, we find that 58.4% of the theorems are proved within 10 minutes, 73.4% within 30 minutes, and 80.8% within 2 hours. Therefore, the majority of proofs are obtained within minutes to tens of minutes—substantially faster than manual proof development, which often requires hours to days per theorem for complex system properties.
@@ 4.3.3 RQ3: Generalizability to Other Projects
A major concern with training-based methods is their generalizability, namely, whether the approach is only applicable to a specific domain due to potential data leakage or overfitting.
To investigate possible leakage, given that the seL4 proofs are publicly available and that the pre-trained model may have already incorporated some of the knowledge, we compute the similarity between the ground-truth proofs and the proofs generated by our approach using sequence similarity (levenshtein1966binary) and Jaccard similarity (jaccard1901etude). As shown in Figure 6, both metrics remain low and, crucially, decrease as proof length grows. If the model were memorizing ground-truth proofs, we would expect similarity to remain high regardless of length; the observed downward trend instead suggests the model is synthesizing novel proof strategies rather than recalling memorized ones. The only exceptions are one- or two-line proofs, where the solution space is inherently small—yet even among these, only 16.7% are identical to the ground truth.
Figure 6: Sequence and Jaccard similarity between generated and ground-truth proofs across proof lengths. Both metrics remain low and decrease as proofs grow longer, suggesting the model synthesizes novel strategies rather than recalling memorized proofs.
We further explore the generalizability of our approach and the fine-tuned LLM. We select three projects from the Archive of Formal Proofs (AFP***https://www.isa-afp.org/), i.e., X86 Semantics X86_Semantics-AFP, IEEE Floating Point IEEE_Floating_Point-AFP, and SATSolverVerification SATSolverVerification-AFP, to directly apply proof search on the theorems therein. In addition, we include the code verification benchmark Code2Inv and translate its loop invariants into Isabelle theorems.
@@ 4.3.4 RQ4: Ablation study
Since our approach integrates several components, namely LLMs, hammer, as well as proof-step revision and proof-state filtering mechanisms, we conduct ablation studies to evaluate the contribution of each part of the framework. To assess the impact of our fine-tuned model, we replace it with off-the-shelf LLMs such as GPT-4o or DeepSeek. Due to budget constraints (running tree search costs approximately \$4.1 per theorem on average when using the DeepSeek or GPT-4o APIs), we randomly sample 200 theorems from the benchmark and use the DeepSeek API to propose proof steps.
In addition, inspired by Rango (thompson2024rango), we implement a retrieval-augmented generation (RAG) mechanism. The retrieval module fetches both similar proof steps and potentially relevant premises. Similar steps are retrieved from the training set based on the similarity between their associated proof states and the current proof state, where these states are tokenized into sequences of identifiers, and relevance is measured using BM25 (robertson2009probabilistic). For premises retrieval, we continue to rely on Sledgehammer’s MePo heuristic for efficiency. The final prompt for LLMs includes the retrieved steps and premises along with their statements, and we restrict the total lengths of steps and premises to 1024 and 512 tokens, respectively.
@@
We also evaluate the impact of the fine-tuning process by adopting a hammer-free tree search setting to focus on the quality of proof steps generated by the model itself. As shown in Table 4, vanilla base models such as Mistral-7B and Qwen3-1.7B struggle with formal proof tasks, yielding negligible success rates even when augmented with tree search. This underscores the necessity of our domain-specific fine-tuning.
@@
Regarding the necessity of proof state filtering, particularly the counterexample detection, our experiments show that incorporating this component allows the system to identify an average of 1.3 counterexamples per evaluated theorem. For the subset of search traces that contain counterexamples, the average number rises to 11.1. These results indicate that counterexample detection is valuable for identifying unprovable states early and preventing unproductive search efforts, although it does not directly increase the overall success rate.
@@
Table 4: Effect of supervised fine-tuning under the same tree-search configuration. Without fine-tuning, the base models solve almost no theorems; in contrast, after supervised fine-tuning, both models reach around 60% proof success.
Method Variation Rate
Mistral-7B + Tree Search 0.0
Qwen3-1.7B + Tree Search 0.1
Mistral-7B SFT + Tree Search 59.8
Qwen3-1.7B SFT + Tree Search 57.1
5 Further Analysis
Beyond aggregate accuracy, we examine how our proposed framework actually constructs successful proofs and what causes the remaining failures.
5.1 Successful Case Studies
We discuss two successful proofs that exhibit complementary patterns: a short symmetry lemma closed by an LLM-proposed case analysis combined with the symbolic tool Sledgehammer (C1), and a long Hoare-triple proof for an seL4 fault handler built entirely from LLM-generated steps (C2). For each, we show the theorem, the proof our framework discovered, and the search behavior behind it.
C1. A symmetry lemma in CSpace. The seL4 lemma same_object_as_commute, from the CSpace invariant development, states that the capability-equivalence relation same_object_as is symmetric. The seL4 proof reduces the equality to a one-way implication via subgoal_tac and closes both directions with a single auto call. Our framework takes a different route: it splits on c first and lets Sledgehammer discharge each constructor case independently.
What the framework did. At depth 0, the LLM proposed several decompositions. Tree search explored seven of them, including case_tac c with various simp-set expansions, before settling on case_tac c followed by simp_all (split: cap.splits), which left a small frontier of constructor cases. Sledgehammer then closed the leaves with single-line blast, meson, and force calls, together with the previously proved same_aobject_as_commute. C2. Invariance under VM-fault handling. The lemma hv_inv_ex, from the ARM ArchSyscall_AI session, is a top-level Hoare triple for seL4’s VM-fault handler: it asserts that an arbitrary precondition $P$ is preserved along the exceptional return path. The seL4 proof is a single composite tactic that relies on dedicated *_inv weakest-precondition lemmas: Our framework arrives at a substantially different proof. It does not use any of the *_inv lemmas; instead, it rebuilds the argument as a 29-step script of wp/wpsimp calls, clarsimp unfoldings, and four explicit postcondition weakenings via hoare_post_imp. Sledgehammer is never invoked: every step is produced by the LLM and validated by the Isabelle kernel. The proof below is abridged. What the framework did. After the initial cases vp split, the proof unfolds along three near-parallel branches, one per fault sub-handler (getDFSR, getFAR, getIFSR). Tree search repeatedly favored wp/wpsimp steps interleaved with clarsimp unfoldings as the symbolic state became exposed. At four points, the LLM proposed intermediate postconditions through rule_tac P=... + hoare_post_imp, which were essential to keep the proof state tractable. None of the 29 steps required Sledgehammer. Summary of successful cases. In both cases, the framework reaches the theorem through a proof structure that differs from the seL4 script: decomposing the equality on the opposite side in C1, and expanding a one-line composite tactic into a 29-step derivation in C2. These examples show that the system genuinely composes LLM generation, symbolic validation, and Sledgehammer integration on non-trivial seL4 goals, rather than replaying memorized scripts. Further examples are available in our GitHub repository.
5.2 Failure Analysis
We next turn to the failures. Across the three evaluation splits, 692 theorems remain unsolved (validation: 241, test: 199, test-hard: 252). We study them at two different granularities. At the step level, our tactic-error analysis draws on the 7.77M individual tactic attempts logged for the 241 failed theorems on the validation split, where each rejected step records the underlying Isabelle kernel error. At the run level, our search-outcome analysis draws on the 451 failed runs from the test and test-hard splits, whose logs record the maximum search depth reached before the budget was exhausted.
Step-level error categories. Of the 7.77M tactic attempts on the validation failures, 30.1% advance or duplicate the proof state, while 69.8% raise a kernel error. Table 5 groups these errors into 12 categories, two of which dominate. The first is lemma hallucination, which combines the undefined-fact, undefined-method, and undefined-constant categories and accounts for 61.1% of all errors. The second is shape mismatch, captured by tactic-inapplicable, which contributes another 17.9%. Together, these two patterns explain 79% of rejected tactics: the model most often cites premises absent from the active context, and otherwise produces well-formed tactics that do not fit the current goal shape. The remaining categories, including schematic-variable mistakes, syntax errors, type errors, tactic timeouts, and unfinished by closures, each contribute under 3%.
Table 5: Step-level error breakdown over all tactic attempts on the validation failures (241 theorems, 7.77M attempts). Undefined-fact alone accounts for nearly two-thirds of all errors: the LLM repeatedly cites premises that do not exist in the active context.
Error category Count % Error category Count %
Undefined fact 3,263,494 60.1 Bad method args. 65,430 1.2
Tactic inapplicable 976,398 17.9 Inner syntax error 50,033 0.9
Multi-line / other 423,982 7.8 Undefined method 47,163 0.8
Unknown variable 161,842 2.9 Proof unfinished 37,100 0.6
Outer syntax error 153,037 2.8 THM exc. 26,750 0.4
Type unification 110,717 2.0 Undefined constant 8,852 0.1
Tactic timeout 102,099 1.8
All errors 5,427,419 100
Search-level outcomes. Table 6 shows how far search gets before exhausting its budget. Only 3.7% of failures stall by depth 10. Most failures therefore make real multi-step progress ($\geq 11$ depth in 96% of cases) but run out of the 128-attempt budget later in the search. On the validation split, where QuickCheck signals are logged, 23 of 241 failed theorems (9.5%) contain at least one search node with a genuine counterexample. These cases suggest that the extracted obligation is unprovable in the active context, typically because an essential assumption is dropped when proof states are sliced from the original seL4 scripts to form the benchmark; we treat them as benchmark-side issues rather than framework failures. Sledgehammer was invoked on all 241 validation failures that still had a residual goal but closed none of them, indicating that the remaining goals lie outside the reach of off-the-shelf SMT/superposition.
Table 6: Search-depth distribution for failures on the test and test-hard splits (the validation log format does not record depth markers). Only 3.7% of failures stall before depth 10; most failures make substantial multi-step progress and run out of budget later in the search.
Max search depth reached Theorems %
0 (no progress past initial state) 11 2.4
111010 (early stall) 6 1.3
11113030 (medium-depth stall) 109 24.2
31316060 (deep stall) 181 40.1
>60>60 (very deep) 145 32.0
Total 451 100
Summary of failure modes. The two analyses converge on the same bottlenecks. At the step level, wasted budget is dominated by missing premises ($\sim$61% of errors) and goal-shape mismatches ($\sim$18%). At the run level, most failed runs do not stall early; they reach substantial depth and then exhaust the budget on residual goals that Sledgehammer cannot close either. We therefore see two clear directions for future work: better premise grounding to reduce lemma hallucination, and longer-horizon planning to prevent late-search collapse on deep proofs.